OpenStudy (anonymous):

What is the solution set for |–4x – 4| – 28 = –4? I got how to do it but everytime I do it fpr the second answer I always mess up

OpenStudy (asnaseer):

can you please show your working so I can spot where you might be making a mistake?

OpenStudy (anonymous):

-4x-4-28=-4 -4x-32=-4 -4x=28 -7 = 1st of the 2 4x-4-28=4 4x-32=4 4x=36 9= second of the 2 but here are my choices a.x = –7 and x = 5 b. x = –7 and x = –5 c. x = 7 and x = 5 d. x = 7 and x = –5

OpenStudy (asnaseer):

ok, it is best to first simplify the equation. so we start with: |-4x - 4| - 28 = -4 |-4x - 4| = -4 + 28 = 24 (adding 28 to both sides) |-x - 1| = 6 (dividing both sides by 4)

OpenStudy (zepp):

I'd suggest to isolate the stuffs in | |

OpenStudy (asnaseer):

now take the two case: a) when x is positive b) when x is negative

OpenStudy (asnaseer):

Case (a) x is positive ============= |-x - 1| = 6 |-(x + 1)| = 6 x + 1 = 6 x = 5

OpenStudy (asnaseer):

do you follow this so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So it has to be a because one is negative and one is positive

OpenStudy (asnaseer):

Case (b) x is negative ============= |-x - 1| = 6 |y - 1| = 6 (let y = -x which makes y a positive number) y - 1 = 6 y = 7 x = -y = -7

OpenStudy (anonymous):

It is A then:D

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Mind if you help me with a little more questions?

OpenStudy (asnaseer):

just post your questions to the list on the left and people will come on and help you.

OpenStudy (asnaseer):

I'll help as well if I'm free :)